# Nonlinear phase synchronization and the role of spacing in shell models

### v1

@@ III.1 Shell Spacing Dependence
Although the widely adopted choice in literature, is to set the inter-shell spacing $g=2,$ having actual triadic interactions involving a Fourier grid structure with the wave-numbers given by $k_{n}$ requires an inter-shell ratio smaller than the golden mean. It is therefore interesting to study how the intermittency corrections depend on the inter-shell-spacing $g$. In a recent work [36], exploring a class of shell models based on recurrent sequences, where the shell wave numbers $k_{n}$ are integers, we observed that the intermittency corrections increase for sequences characterized by smaller inter-shell spacings.
@@
Figure 5: Plot of the Scaling exponent $\xi_{p}$of the $p-$th order structure function as a function of $p$, for different shell spacings considered. The dashed line represents the Kolmogorov $p/3$ scaling.
@@
Figure 6: Scaling exponent $\xi_{p}$for structure functions of order $p$ (ranging from 4 to 10), plotted as a function of the inter-shell spacing $g$. Colors of the markers correspond different shell spacings as in Fig. 5.
@@
In order to study this phenomenon, here we provide a detailed characterization of the intermittency as a function of inter-shell spacing, in particular considering the role of triadic synchronization in the generation of rare and intense events in the energy transfer responsible for the increased intermittency.
@@
We first present how intermittency corrections display a non-trivial dependence on the inter-shell spacing $g$ for the standard (GOY) shell model in Figure 5, which shows the estimated scaling exponents $\xi(p,g)$ of the $p$-th order structure functions, for different values of the spacing, in the range $g=1.1-2$, revealing a progressive enhancement of intermittency, corresponding to a flattening of the curves of scaling exponents in the small $g$ limit. We note that those observations are consistent with previous studies, especially investigating the continuum limit of shell models, where the $g\to 1$ limit leads to equations closely resembling the Burgers equation [55], which is well known to generate shocks and an intermittent bifractal anomalous scaling behavior [56]. Further insight is provided by Figure 6, which shows the dependence of the scaling exponent $\xi(p,g)$ as a function of the inter shell spacing $g$. Note how all the scaling exponents exhibit a sharp variation when $g$ decreases below a critical value $g_{c}\simeq 1.7$, indicating a transition in the scaling properties of the system. For $g>g_{c}$ there is no relevant modification with respect the inter shell spacing, while for $g<g_{c}$ the anomalous exponents vary significantly with respect to the inter shell spacing, following an approximately linear scaling.
@@
Note also how $g_{c}$ is just slightly above the golden ratio. While this may be a pure coincidence, there may be a link to nonlinear frequencies being close enough to interact more efficiently through recurrence relations such as $\omega_{n}=\omega_{n-1}+\omega_{n-2}$.
@@ III.1.1 Shell spacing and the level of synchronization
With the aim of highlighting the role of synchronization events in enhancing intermittency corrections, we analyze the statistical properties of triadic phases across the inertial range. In figure 4, we present the scaling of different structure functions for two representative cases, corresponding to $g=2$ and $g=1.4$, respectively characterized by a weak (or usual) and a strong degree of intermittency. In the inset of the same figure, we present the PDFs of triadic phases for different shells within the inertial range. In the case of higher intermittency (smaller $g$) the phase distributions display a clearly scale-dependent behavior, becoming progressively narrower for smaller scales in the inertial rage. This suggests a tendency towards a stronger phase alignment for higher wave numbers.
In order to characterize the level of synchronization at a global level across the inertial range, we consider the complex Kuramoto order parameter as defined earlier in Sec. II.3 and present its PDFs in Fig. 7 as a function of the inter-shell spacing $g$.
We note that, as the inter-shell spacing decreases, the PDFs become increasingly skewed and a finite probability for fully synchronized events appears. These events correspond to states in which all the triads are aligned, and are associated with rare, extreme events in the energy flux. Remarkably, we observe that below the critical value $g<g_{c}$, the PDFs of $R$ are $g$-dependent, while for for $g>g_{c}$ they collapse on the same distribution, which is also in accordance with the behavior of intermittency corrections as a function of $g$. This suggests that the Kuramoto parameter, despite being only a measure of the phase coherence in the systems, can be used as a proxy for detecting different levels of intermittency.
The statistics of the average phase, represented by the phase $\Phi$ of the Kuramoto parameter is presented in Fig. 8. The PDFs of $\Phi$ as a function of $g$ are all centered around $\pi/2$, but with a width that decreases rapidly for $g<g_{c}$ , indicating that for smaller $g$ the system has a stronger tendency to cluster around a common phase that maximizes the forward energy flux.
@@
Figure 8: Normalized PDFs of the average phase $\Phi$, for different shell spacings $g$.
Not surprisingly, the emergence of global synchronization is a manifestation of the presence of local coherence. This can be seen clearly in figure 9, which is equivalent to Fig. 3, but with an inter shell spacing of $g=1.5,$ corresponding to strong intermittency. In Panel $(a)$, the statistics the local spectral energy flux $\Pi_{n}^{E}$ are presented, showing an increasing occurrence and intensity of energy transfer along with a reduced probability of negative flux events as we move towards smaller scales. This behavior is evidenced by the progressively heavier tails of the distribution curves. The scale-dependent organization of triadic phases $\varphi_{n}$ is clarified by the statistics of the local phase coherence parameter $R_{n},$ shown in panel $(b)$, which reveal that smaller scales display higher synchronization levels, with an increasing probability of fully coherent states ($R_{n}\simeq 1$). These events correspond to clusters of partially aligned triads with $\varphi_{n}\simeq\pi/2$ near a stable fixed point of the phase dynamics.
The correlation between the energy transfer and the local synchronization level along the triads chain is analyzed via the joint statistics of $\Pi_{n}^{E}$ and $R_{n}$, as well as $\Pi_{n}^{E}$ and $\Phi_{n}$, shown in panel (c) and (d) respectively. These plots demonstrate that highly local synchronized triad configurations - identified by $R_{n}\simeq 1$ and by local phase alignment $\Phi_{n}\simeq\pi/2$ are strongly correlated with extreme flux events.
Figure 9: Individual and joint statistics of energy flux and the amplitude and phase of the Kuramoto parameter. In panel (a) on the top left plot, the PDF of the local spectral energy flux $\Pi_{n}^{E}$ is shown, exhibiting the typical behaviour where the distributions becomes progressively more heavy-tailed for shells corresponding to smaller scales. This reflects the increasing occurrence and intensity of energy bursts as we move towards smaller scales in the inertial range. In panel (b) on the top right, the PDF of the local Kuramoto phase coherence $R_{n}$ is presented, where smaller scales display higher synchronization levels, with an increasing probability of fully coherent local $R_{n}$. Curve colors correspond to different shell numbers, as indicated by the continuous colorbar. In panel (c) on the bottom left, the joint PDF between $\Pi_{n}^{E}$ and $R_{n}$ is shown, which illustrates how extreme flux events are predominantly associated with high local synchronization events. In panel (d) on the bottom right, the joint PDf between $\Pi_{n}^{E}$ and $\Phi_{n}$ is shown, complementary to panel (c), reinforcing the observation that extreme flux events are associated with a marked clustering of triad phases around $\pi/2$. Note that this state, where the phases are locked, is highly efficient in transferring energy towards small scales.
@@ III.2 Burst events and temporal intermittency
So far, we have characterized the statistical properties of shell models as a function of the inter-shell spacing $g$, showing the association between phase coherence and statistics of extreme flux events, ultimately leading to different levels of intermittency.
In particular, we showed the flattening of the scaling exponents $\xi(p)$ of the structure functions, observed for decreasing values of $g$ similar to what happens in the Burgers equation with the formation of shocks. In shell models, this correspond to a dynamics dominated by burst-like solutions, in the form of nonlinear traveling waves that in the continuum limit has been addressed analytically by assuming a zero background by Andersen et al. [55].
These burst events are present also in the fully developed state together with a background power law solution. They can be identified as a sequence of coherent maxima propagating across scales, which are sometimes referred to as instantons, and have been proposed as a key mechanism leading to anomalous scaling in shell models [20, 21].
A detailed statistical characterization of these instantons as a function of the inter-shell spacing $g$ lies beyond the scope of the present work. This would represent an interesting future direction, particularly in the low- $g$ regime, where we observe also backward-propagating events, suggesting a scenario more akin to a gas of colliding structures, rather than isolated, unidirectional bursts.
Nevertheless, continuing with the line of inquiry of this work, we present a phenomenological analysis of these burst like events with respect to the synchronization properties of the system. First we present how the dynamics of these events depend on the degree of intermittency, controlled by the inter-shell spacing $g$.
Before that however, in order to visualize these structures, Fig. 10 provides the temporal evolution of the compensated shell amplitudes, for different values of the inter-shell ratio $g=2,1.4,1.2$. We can see rather frequent but less intense bursts for larger $g$, some of which may dissipate or annihilate within the inertial range, before reaching the smallest scales. In contrast, as $g$ is decreased, bursts become rarer, more localized, and significantly more intense, allowing us to see their propagation across the whole inertial range. In this regime, a given shell spends most of the time in a quiescent state, interrupted by the passage of large-scale coherent bursts, which may also occasionally be accompanied by secondary events forming tree-like patterns.
@@
In order to support our hypothesis that the dynamics of these bursts are associated with phase synchronization, we present the time evolution of compensated amplitudes and energy flux, and the local Kuramoto parameter $R_{n}$ in Figure 11, for inter-shell spacing $g=1.5$. This demonstrates that the approach of a burst corresponds to a rapid growth of local synchronization level, which precedes the extreme flux event. It also shows that as the flux event arrives at the smallest scales, all the shells are synchronized and they loose this synchronization slowly over time with the smallest scales becoming desynchronized first. In this sense, synchronization starts from large scales and propagate towards small scales while desynchronization starts from small scales and propagates towards large scales.
[…]
During this period, energy injected at large scales tends to accumulate again in the low-$n$ shells, and a new burst is eventually triggered, in a sort of cyclic dynamics of energy accumulation and release. (Interesting our observation echos, with the results of Ref. [57] in which avalanche-like behavior of seismic events, has been compared to energy release events and their inter-event times (waiting times). In our framework, the duration of quiescent states — i.e., the time intervals during which energy accumulates at large scales — appears to be directly influenced by the inter-shell spacing $g$.
@@ IV Helical shell models and sync
We have shown that the existence of phase coherent energy transport events account for the anomalous scalings of structure functions and therefore intermittency in shell models, and that the phase coherence happens through a mechanism of nonlinear synchronization. This means that the affinity of its triads to synchronization, is the key feature of a shell model that determine its intermittency. So far, we have characterized the synchronizability of only the simplest shell model: the classical GOY model with a one dimensional chain consisting in $N$ shells connected by $N-2$ triads.
However, the synchronizability of a system of triads (or, generally, a system of coupled oscillators) also depends on the way its elements are wired, which is represented by its topology. It is somewhat obvious that topological effects became relevant as we depart from a linear chain picture, as in the case of more complex model such as nested polyhedra models or log-lattices, where multiple elements interact at each scale, but even if we put aside those more complex examples, a minimal reintroduction of network topology appears in simple shell models if we try to incorporate helicity in a proper way. This happens if we consider two signs of helicity in each scale (instead of one sign in odd scales and another sign in even scales) as is the case in helical shell models, a natural generalization of the standard shell model, in which each shell is assigned two dynamical variables $u_{n}^{+},$$u_{n}^{-}$ carrying respectively positive and negative helicities. In this framework the two inviscid invariants are defined as:
@@
This formulation results in two, one-dimensional sequences of shells variables, corresponding to the two helicity signs. The triadic interactions among those modes can be systematically recovered from Waleffe’s helical decomposition [58], where the nonlinear term in the Navier-Stokes equations is decomposed into distinct classes of helical interactions.
The general form of a helical shell model, omitting the dissipative and forcing terms, with arbitrary non-local triads connecting the three shells $n-l,n,n+m$, can be written compactly as:
@@
$s_{3}\equiv s_{0}s_{1}s_{2},$ and where $Q=Q_{\ell m}\left(s_{1}/s_{0},s_{2}/s_{0}\right)$ is the geometric factor of the triad class that is considered. The four triad classes can be defined as the $4$ possible combinations of the signs of $s_{1}/s_{0}$ and $s_{2}/s_{0}.$ Notice that the choice of local interactions corresponds to setting $l=1,m=1$, and results in $2N$ modes connected by $42N-$ 4 triads for each class. This represents - with the usual choice $g=2$- the helical shell model introduced in Ref. [59], investigated by considering the different classes of interactions separately.
However, the general formulation of the model allows to consider arbitrary elongated triads as investigated by De Pietro, Biferale, and Mailybaev [60]. Moreover, one can also consider a right-hand side that constitutes a sum over different classes, each corresponding to different geometrical factors defined in terms of the shell index and the elongation of the triads [40].
Each of these choices correspond to a different network of triadic interactions, connecting two sequences of helical modes. For simplicity, we focus on models constructed from only a single interaction class, and adopt the labeling SM1, SM2, etc. (see , e.g. , Ref. [59, 60]) following Benzi and De Pietro. There are qualitative differences among the triadic networks of different classes, in particular, for certain classes of models, the triadic interactions form two independent chains that do not interweave. We call those models separable.
This allows us to distinguish two types of topologies namely separable and non separable. The SM4 class, corresponds to the configuration $s_{1}/s_{0}=$$s_{2}/s_{0}=+$, is separable since it consists of homochiral triads of the type (+,+,+)(+,+,+)( + , + , + ) or (,,)(-,-,-)( - , - , - ) where two helicities decouple and each helical variable evolves independently that conserve separately their corresponding positive-definite invariants, (i.e. $\sum_{n}k_{n}\left|u_{n}^{+}\right|^{2}$and $\sum_{n}k_{n}\left|u_{n}^{-}\right|^{2}$), making it a candidate to drive an inverse cascade, which is in practice overwhelmed by equipartition. The other separable class is the SM1 class, corresponding to $s_{1}/s_{0}=-$, and $s_{2}/s_{0}=-,$ involving heterochiral triads of the type (,+,)(-,+,-)( - , + , - ) or (+,,+)(+,-,+)( + , - , + ), which naturally splits into two distinct chains of modes with alternating helicities, consisting effectively of two uncoupled GOY models.
@@
In this case, the right-hand side of the phase evolution equation involves contributions from up to seven different triads, in contrast to the five involved in the separable (single-chain) case, as detailed in Sec. II.1. Therefore, in order to characterize the level of synchronization in a non separable chain, we define a weighted local Kuramoto parameter, constructed by summing the contributions of up to seven phases corresponding to the interacting triads.
@@
$R_{n}e^{i\Phi_{n}}\equiv\frac{\sum_{t\in\Gamma(n)}w_{n}e^{i\varphi_{n}}+\bar{w}_{n}e^{-i\bar{\varphi}_{n}}}{\sum_{t\in\Gamma(n)}w_{n}+\bar{w}_{n}}\;\text{,}$ (9)
@@
are derived from the structure of the phase evolution equation that correspond to set of interacting triads $\Gamma(n)$ which are connected with the $n$-th shell triad $u_{n-1}^{s_{1}}u_{n}^{s_{0}}u_{n+1}^{s_{2}}=\chi_{n}e^{i\varphi_{n}}.$ These weights depends on the topology of each class, e.g., for the triads of type SM3 (+,+,)(+,+,-)( + , + , - ) the weights are $\left\{0,1,3,1,0;1,0,0,1,1\right\}$, while for the separable class SM1 (,+,)(-,+,-)( - , + , - ) they are $\left\{1,0,3,0,1;0,2,0,2,0\right\}$ consistently with the definition 7 for the single chain case.
In the following, we will characterize the level of synchronization in helical shell models according to their cascade properties and the presence of intermittency corrections.

### v2

@@ III.1 Shell Spacing Dependence
Although the widely adopted choice in literature, is to set the inter-shell spacing $g=2,$ having actual triadic interactions involving a Fourier grid structure with the wave-numbers given by $k_{n}$ requires an inter-shell ratio smaller than the golden ratio. It is therefore interesting to study how the intermittency corrections depend on the inter-shell spacing $g$. In a recent work [36], exploring a class of asymptotically self-similar shell models based on recurrent sequences -where the shell wave numbers $k_{n}$ are integers- we observed that the intermittency corrections increase for sequences characterized by smaller inter-shell spacings.
@@
Figure 5: The scaling exponent $\xi(p)$ of the $p$-th order structure function as a function of $p$, for different shell spacings. The dashed line represents the Kolmogorov $p/3$ scaling.
@@
Figure 6: Scaling exponent $\xi_{p}(g)$ for structure functions of order $p$ (ranging from 4 to $10$ in steps of $0.5$), plotted as a function of the inter-shell spacing $g$. The marker colors correspond to different shell spacings, as in Fig. 5.
@@
In order to study this phenomenon, here we provide a detailed characterization of the intermittency as a function of inter-shell spacing in regular shell models, in particular considering the role of triadic phase synchronization in the generation of rare and intense events of energy transfer that are responsible for the increased intermittency.
@@
We first note that the intermittency corrections have a non-trivial dependence to the inter-shell spacing $g$ for the standard (GOY) shell model. Figure 5 shows the estimated scaling exponents $\xi(p,g)$ of the $p$-th order structure functions, for different values of the spacing, in the range g=1.12g=1.1-2, suggesting a progressive enhancement of intermittency, corresponding to a flattening of the scaling exponent curves as $g$ is decreased. These observations are consistent with previous studies, particularly those investigating the continuum limit of shell models, where $g\to 1$ leads to equations closely resembling the Burgers equation [56], known to generate shocks and an intermittent, bifractal anomalous scaling behavior [57]. Further insight is provided by Figure 6, which shows the dependence of the scaling exponents $\xi_{p}(g)$ as a function of the inter-shell spacing $g$. High order scaling exponents such as $p\gtrsim 8$ exhibit a sharp transition when $g$ decreases below a critical value $g_{c}\simeq 1.7$, suggesting a qualitative change in the scaling properties of the system. It seems that for $g\apprge g_{c}$ the scaling is roughly independent of the inter-shell spacing, while for $g\apprle g_{c}$ the anomalous scaling exponents vary significantly with the inter-shell spacing, following an approximately linear scaling.
@@
Note also how $g_{c}$ is just slightly above the golden ratio. While this may be a pure coincidence, there may be a link to nonlinear frequencies $\omega_{n}$, of the $n$-th triad being close enough to interact more efficiently through a recurrence relation such as $\omega_{n}=\omega_{n-1}+\omega_{n-2}$.
@@ III.1.1 Shell spacing and the level of synchronization
With the aim of highlighting the role of synchronization events in increasing intermittency corrections, we analyze the statistical properties of triadic phases across the inertial range. In Figure 4, we present the scaling of structure functions for two representative cases, $g=2$ and $g=1.4$, which are characterized by a standard and very high levels of intermittency, respectively. In the inset of the same figure, we present the PDFs of triadic phases for different shells within the inertial range. For the higher intermittency case ($g=1.4$), the triadic phase distributions exhibit a pronounced scale-dependent behavior, becoming progressively narrower at smaller scales in the inertial range, a similar trend is also observed in the statistics of the Kolmogorov multipliers (not shown in the plot). This behavior is actually common for all the $g$ values below the critical value $g\apprle g_{c}$ and becomes more pronounced as $g$ decreases. This suggests a tendency towards stronger phase alignment at higher wave numbers, consistent with the idea that the increase in intermittency is due to rare but large dissipation events at small scales.
In order to characterize the global level of synchronization across the inertial range, we consider the complex Kuramoto order parameter, defined in Eq. (7), and present the PDFs of the coherence parameter $R$ in Fig. 7 as a function of the inter-shell spacing $g$.
We note that, as the inter-shell spacing decreases, the PDF of the coherence parameter $R$ become increasingly skewed and a finite probability for fully synchronized events appears. These are states in which all the triads are aligned, and are associated with rare, extreme events in the forward energy flux. Remarkably, we observe that below the critical value $g\apprle g_{c}$, the PDFs of $R$ are $g$-dependent, while for for $g\apprge g_{c}$ they collapse to the same distribution, which is also in accordance with the behavior of intermittency corrections as a function of $g$. This suggests that the Kuramoto parameter, despite being only a measure of the phase coherence, can be used as a proxy for detecting different levels of intermittency.
The statistics of the average phase, represented by the phase $\Phi$ of the Kuramoto parameter is presented in Fig. 8. The PDFs of $\Phi$ as a function of $g$ are all centered around $\pi/2$, but with a width that decreases rapidly for $g\apprle g_{c}$ , indicating that for smaller $g$ the system has a stronger tendency to cluster around a common phase that maximizes the forward energy flux.
@@
Figure 8: Normalized PDFs of the average phase $\Phi$, for different shell spacings $g$. Marker colors correspond to the different shell spacings, as in Fig. 7.
Not surprisingly, global synchronization goes hand in hand with local coherence. This can be seen clearly in Figure 9, which is equivalent to Fig. 3, but with an inter-shell spacing of $g=1.5,$ corresponding to strong intermittency. In Panel $(a)$, the statistics of the local spectral energy flux $\Pi_{n}^{E}$ are presented, showing an increasing occurrence and intensity of energy transfer along with a reduced probability of negative flux events as we move towards smaller scales. This behavior is evidenced by the progressively heavier tails of the distribution curves. The scale-dependent organization of triadic phases $\varphi_{n}$ can be seen in the statistics of the local phase coherence parameter $R_{n}$, in panel $(b)$, which show that smaller scales are associated with higher synchronization levels, and an increased probability of fully coherent states ($R_{n}\simeq 1$). These events are clusters of locally aligned triads with $\varphi_{n}\simeq\pi/2$ corresponding to a fixed point of the phase dynamics.
The correlation between the energy transfer and the local synchronization level along the triad chain is analyzed via the joint statistics of $\Pi_{n}^{E}$ and $R_{n}$, as well as $\Pi_{n}^{E}$ and $\Phi_{n}$, shown in panel (c) and (d) respectively. These plots demonstrate that local synchronized triad configurations - identified by $R_{n}\simeq 1$ and $\Phi_{n}\simeq\pi/2$ - are strongly correlated with extreme flux events.
Figure 9: Individual and joint statistics of energy flux and the amplitude and phase of the Kuramoto parameter. In panel (a) on the top left plot, the PDF of the local spectral energy flux $\Pi_{n}^{E}$ is shown, exhibiting the typical behavior where the distributions becomes progressively more heavy-tailed for shells corresponding to smaller scales. This reflects the increasing occurrence and intensity of energy bursts as we move towards smaller scales in the inertial range. In panel (b) on the top right, the PDF of the local Kuramoto phase coherence $R_{n}$ is presented, where smaller scales display higher synchronization levels, with an increasing probability of local fully coherent states ($R_{n}\simeq 1$). Curve colors correspond to different shell numbers, as indicated by the continuous colorbar. In panel (c) on the bottom left, the joint PDF between $\Pi_{n}^{E}$ and $R_{n}$ is shown, which illustrates how extreme flux events are predominantly associated with high local synchronization events. In panel (d) on the bottom right, the joint PDf between $\Pi_{n}^{E}$ and $\Phi_{n}$ is shown, complementary to panel (c), reinforcing the observation that extreme flux events are associated with a marked clustering of triad phases around $\pi/2$. Note that this state, where the phases are locked, is highly efficient in transferring energy towards small scales.
@@ III.2 Burst events and temporal intermittency
So far, we have characterized the statistical properties of shell models as a function of the inter-shell spacing $g$, showing the relation between phase coherence and statistics of extreme flux events, leading to different levels of intermittency.
In particular, we showed that the flattening of the scaling exponents $\xi(p)$ of the structure functions that is observed for decreasing values of $g$, is rather similar to what happens in Burgers equation with the formation of shocks. The equivalent process in shell models is the formation of burst-like solutions. These solutions have been addressed in the continuum limit analytically by assuming a zero background by Andersen et al. [56].
Similar burst events are also present in the fully developed turbulent state, coexisting with a background power law solution of the average spectral energy. They can be identified as coherent sequences of local maxima in the time evolution of the shell amplitudes $U_{n}$, propagating across shells. Such structures, sometimes referred to as instantons, and have been proposed as a key mechanism underlying anomalous scaling in shell models [20, 21].
A detailed characterization of the statistical properties of instantons as a function of the inter-shell spacing $g$ lies beyond the scope of the present work. This would represent an interesting future direction, particularly in the low-$g$ regime, characterized by stronger intermittency, where we observe also backward-propagation events, suggesting a scenario along the lines of a gas of colliding structures, rather than isolated, unidirectional bursts.
Nevertheless, continuing with the line of inquiry of this work, we present a phenomenological analysis of these burst like events with respect to the local synchronization properties of the system.
First we present how the dynamics of these events depend on the degree of intermittency, controlled by the inter-shell spacing $g$. In order to visualize these structures, Fig. 10 provides the temporal evolution of the compensated shell amplitudes, for different values of the inter-shell ratio $g=2,1.4,1.2$. We can see rather frequent but less intense bursts for larger $g$, some of which may dissipate or annihilate within the inertial range, before reaching the smallest scales. In contrast, as $g$ is decreased, bursts become less frequent, more localized, and significantly more intense, allowing us to see their propagation across the whole inertial range. In this regime, a given shell spends most of its time in a quiescent state, interrupted by the passage of large-scale coherent bursts, which may also occasionally be accompanied by secondary events forming tree-like patterns.
@@
In order to support our hypothesis that the dynamics of these bursts are associated with triadic phase synchronization, we also present the time evolution of the compensated amplitudes, the energy flux, and the local Kuramoto parameter $R_{n}$ in Figure 11, for inter-shell spacing $g=1.5$. This demonstrates that the arrival of a burst corresponds to a rapid growth of local synchronization level, which precedes the extreme flux event. It also shows that as the flux event arrives at the smallest scales, all the shells are synchronized and they loose this synchronization slowly over time with the smallest scales becoming desynchronized first. In this sense, synchronization starts from large scales and propagates towards small scales while desynchronization starts from small scales and propagates towards large scales.
One explanation of these observations is that the energy, being injected at large scales, tends to accumulate in lown-n shells, until the triadic phases can organize coherently, facilitating the release of this accumulated energy across the shells. This naturally starts at the large scales, and proceeds like a domino effect towards small scales. Therefore, as the burst propagates, the local Kuramoto parameter exhibits a sharp increase, corresponding to a local synchronization of the system. This is necessary to accommodate the passage of the large deviation in flux. After the passage of the burst, the system remains partially phase locked for some time, corresponding to a cluster of scales with $R_{n}\simeq 1$, maintaining an “open channel” for energy transfer (in terms of triadic phases). This synchronized configuration corresponds to a fixed point of the triadic phase equation, which provides a quasi-stationary state. It then gets desynchronized at the smallest scale first, probably because of the nonlinear time scales associated with the smallest scales are the shortest, so the decay to a desynchronized state in the absence of a synchronizing force is faster at small scales. Although we do not present the analogue of Fig. 11 for different values of the inter-shell spacing, we note that as $g$ decreases, the phase locked state appears to be more robust and stable, requiring a longer time for the system to desynchronize and return to a disordered phase configuration.
After the decay of the burst, energy injected at large scales starts to accumulate again in the low-$n$ shells, and a new burst is eventually triggered, in a sort of cyclic dynamics of energy accumulation and release. (Interesting our observation echos, with the results of Ref. 58 in which avalanche-like behavior of seismic events, has been compared to energy release events and their inter-event times (waiting times). In our framework, the duration of the quiescent states — i.e., the time intervals during which energy accumulates at large scales — appears to be directly influenced by the inter-shell spacing $g$.
@@ IV Helical shell models and sync
We have shown that the existence of phase coherent energy transport events account for the anomalous scalings of structure functions and therefore intermittency in shell models, and that the phase coherence happens through a mechanism of nonlinear synchronization. This means that the affinity of its triads to synchronization, is the key feature of a shell model that determine its intermittency. So far, we have characterized the synchronizability of only the simplest shell model: the classical GOY model with a one dimensional chain consisting in $N$ shells connected by N2N-2 triads.
However, the synchronizability of a system of triads (or, generally, a system of coupled oscillators) also depends on the way its elements are wired, which is represented by its topology. It is somewhat obvious that topological effects become relevant as we depart from a linear chain picture, as in the case of more complex model such as nested polyhedra models or log-lattices, where multiple elements interact at each scale, but even if we put aside those more complex examples, a minimal reintroduction of network topology appears in simple shell models if we try to incorporate helicity in a proper way. This happens if we consider two signs of helicity in each scale -instead of one sign in odd scales and another sign in even scales- as in helical shell models. These models are a natural generalization of the standard shell model, in which each scale is associated with two dynamical variables, $u_{n}^{+}$ and $u_{n}^{-}$, carrying positive and negative helicities, respectively. In this framework the two inviscid invariants are defined as:
@@
This formulation results in two, one-dimensional sequences of shell variables, corresponding to the two helicity signs. The triadic interactions among those modes can be systematically recovered from Waleffe’s helical decomposition [59], where the nonlinear term in the Navier-Stokes equations is decomposed into distinct classes of helical interactions.
The general form of a helical shell model, omitting the dissipative and forcing terms, with arbitrary non-local triads connecting the three shells nl,n,n+mn-l,n,n+m, can be written compactly as:
@@
$s_{3}\equiv s_{0}s_{1}s_{2},$ and where $Q=Q_{\ell m}\left(s_{1}/s_{0},s_{2}/s_{0}\right)$ is the geometric factor of the triad class that is considered. The four triad classes can be defined as the $4$ possible combinations of the signs of $s_{1}/s_{0}$ and $s_{2}/s_{0}.$ Notice that the choice of local interactions corresponds to setting $l=1,m=1$, and results in $2N$ modes connected by 2N42N-4 triads for each class. This represents - with the usual choice $g=2$ - the helical shell model introduced in Ref. 60, investigated by considering the different classes of interactions separately.
However, the general formulation of the model allows us to consider arbitrarily elongated triads as investigated by De Pietro, Biferale, and Mailybaev [61]. Moreover, one can also consider a right-hand side that constitutes a sum over different classes, each corresponding to different geometrical factors defined in terms of the shell index and the elongation of the triads [41].
Each of these choices correspond to a different network of triadic interactions, connecting two sequences of helical modes. For simplicity, we focus on models constructed from a single interaction class, and adopt the labeling SM1, SM2, etc. (see , e.g. , Refs. 60, 61) following Benzi and De Pietro. There are qualitative differences among the networks represented by different classes. Note in particular that for certain models, the triadic interactions form two independent chains that do not interweave. We call those models separable.
This allows us to distinguish two types of topologies namely separable and non separable. The SM4 class, corresponds to $s_{1}/s_{0}=$$s_{2}/s_{0}=+$, is separable since it consists of homochiral triads of the type $(+,+,+)$ or (,,)(-,-,-) where two helicities decouple and each helical variable evolves independently that conserve separately their corresponding positive-definite invariants, (i.e. $\sum_{n}k_{n}\left|u_{n}^{+}\right|^{2}$and $\sum_{n}k_{n}\left|u_{n}^{-}\right|^{2}$), making it a candidate to drive an inverse cascade, which is in practice overwhelmed by equipartition. The other separable class is the SM1 class, corresponding to $s_{1}/s_{0}=$-, and $s_{2}/s_{0}=-,$ involving heterochiral triads of the type (,+,)(-,+,-) or (+,,+)(+,-,+), which naturally splits into two distinct chains of modes with alternating helicities, consisting effectively of two uncoupled GOY models.
@@
In this case, the right-hand side of the phase evolution equation involves contributions from up to seven different triads, in contrast to the five involved in the separable (single-chain) case, as detailed in Sec. II.1. Therefore, in order to characterize the level of synchronization in a non separable chain, we define a weighted local Kuramoto parameter, constructed by summing the contributions of up to seven phases of its interacting triads.
@@
$R_{n}e^{i\Phi_{n}}\equiv\frac{\sum_{t\in\Gamma(n)}w_{t}e^{i\varphi_{t}}+\bar{w}_{t}e^{-i\bar{\varphi}_{t}}}{\sum_{t\in\Gamma(n)}w_{t}+\bar{w}_{t}}\;\text{,}$ (10)
@@
are determined by the structure of the phase evolution equation that correspond to a set of interacting triads $\Gamma(n)$ which are connected with the central triad of the $n$-th shell $u_{n-1}^{s_{1}}u_{n}^{s_{0}}u_{n+1}^{s_{2}}=\chi_{n}e^{i\varphi_{n}}.$ These weights, consistently with the previous definition (Eq. (8)), are equal to the multiplicity of each triad in the right-hand side of the evolution of equation for the triadic phase $\varphi_{n}$. In helical shell model, their values depends on the topology of the interaction class. For instance, for the triads of type SM3 the weights are $\left\{0,1,3,1,0;1,1,0,1,1\right\}$, while for the separable class SM1, they are $\left\{1,0,3,0,1;0,2,0,2,0\right\}.$ Please note how these weights are consistent with the definition adopted for the single chain case, that we can express as $\left\{w_{n-2},\bar{w}_{n-1},w_{n},\bar{w}_{n+1},w_{n+2}\right\}$, since the SM1 reduces to two uncoupled GOY models.
In the following, we will characterize the level of synchronization in helical shell models with respect to their cascade properties and the presence of intermittency corrections.